Find the intercepts and asymptotes. (If an answer does not exist, enter DNE. Enter your asymptotes as a comma-separated list of equations if necessary.) r(x) = 5x^2 − 20x + 21 / x^2 − 4x + 4

x-intercept (x, y) =

y-intercept (x, y) =

vertical asymptote(s)

horizontal asymptote
Answer #1

Solution given that equation is 57?- 2031 +21 r(a) = 22 42 + 4 to find X-Intercept uee will put r(1)=y=o So, - 7101) 512 2071For y - Intercept put d = 0 ola) = 51002- 2010) +21 (0)2- 4101 +4 y = 710) alt so, y - Intercept is 10,*) the ule will zero oy = Coeffieclent of numerator of highest degree Corpfiuent of denominator of highest degree her degne of numerator is (2) equ